OpenStudy (anonymous):

Solve x2 + 18 = –9x.

OpenStudy (whpalmer4):

Move all of the terms to the left side by adding \(9x\) to both sides. What do you get?

OpenStudy (anonymous):

x^2 + 9x + 18 = 0 (x+6)(x+3) =0 x = -6,-3

OpenStudy (anonymous):

thats what i got too thanks!